# When Do Transformers Learn Heuristics for Graph Connectivity?

### v1

@@ C.1 Characterizing Block Weights $W_{\ell}$
To do so, the proof is split into two parts: we prove that each $n\times n$ block must be diagonal using (ii) and Section B.2, and that the diagonal entries must realize the said forms by examining the forward maps under a curated, parameterized class of initial hidden states.
Step 1: each block must be diagonal. In this step, we argue that if a block admits a positive off-diagonal entry, then the certificate trick from Section B.2 will create a false positive entry on some output, contradicting (ii).
@@
Consider feeding the graph $G_{0}$ of $n$ isolated vertices into $\mathsf{TF}_{\Theta}^{L}$, so that $h_{0}(G_{0})=[I_{n}\mid I_{n}]$. The premises of Section B.2 hold, so every column of every $h_{\ell}(G_{0})$ is supported in exactly one row, which we called its label in $\{1,\ldots,n\}$. Hence, if we write $h_{\ell}(G_{0})=[X_{1}^{(\ell)}\mid\ldots\mid X_{K_{\ell}}^{(\ell)}]$ of contiguous $n\times n$ blocks, then each such block $X_{r}^{(\ell)}$ must be nonnegative and diagonal. Now expand
@@
  • If both $(i,r)$ and $(k,s)$ are activatable, take graphs $G_{i},G_{k}$ that make $X_{r}^{(\ell-1)}(G_{i})[i,i]>0$ and $X_{x}^{(\ell-1)}(G_{k})[k,k]>0$. Using the certificate mechanism in Section B.2, each positiveness admits a finite certificate subgraph with at most $2\cdot 3^{\ell-1}$ vertices. We then create a new graph $G^{\prime}$ and disjointly embed both certificates into it, leaving all other vertex isolated. The two labels $i,k$, viewed as nodes, now lie in different components. But then the product
@@
Suppose an $L$-layer Disentangled Transformer $\mathsf{TF}_{\Theta}^{L}$ has nonnegative parameters. Suppose $\mathsf{TF}_{\Theta}^{L}$ is layerwise permutation equivariant, i.e., for each $\ell$ and any hidden states $h\in\mathbb{R}^{n\times d_{\ell-1}}$,
@@ C.2 Population Gradient Lives in the Equivariant Algebra
Under Section 4.3, in particular using layerwise parameterization $W_{\ell}=A_{\ell}\otimes I_{n}+B_{\ell}\otimes J_{n}$, fix a layer $\ell$ and let $K=K_{\ell-1}$. Then the population gradient with respect to $W_{\ell}$ lies in $M_{K}(\mathbb{R})\otimes\mathrm{span}\{I_{n},J_{n}\}$: there exist matrices $G_{\ell}^{(I)},G_{\ell}^{(J)}\in\mathbb{R}^{K\times K}$ such that
@@ C.3 Which Conditions Encourage $W_{\ell}\approx A_{\ell}\otimes I_{n}$?
Throughout the analysis of training dynamics, we inherit the notations used in Section 4.3: we use $Z$ to denote the model output, $R$ the reachability matrix, $A$ the adjacency matrix, $\mathcal{L}=\mathcal{L}(Z;R)$ the loss, and $\mathcal{R}(\Theta)$ the population risk $\mathcal{R}(\Theta):=\mathbb{E}_{G\sim\mathrm{ER}(n,p)}[\mathcal{L}(\mathsf{TF}_{\Theta}^{L}(A_{G});R_{G})]$.
@@
Assume Section 4.3. Fix any layer $\ell$ and decompose $W_{\ell}=A_{\ell}\otimes I_{n}+B_{\ell}\otimes J_{n}$. Let $Z$ be the output, $R$ the reachability matrix (ground truth), $\mathcal{L}=\mathcal{L}(Z;R)$ the loss, and $\mathcal{R}(\Theta)$ the population risk.
@@
    $D_{B_{\ell}}\mathcal{R}(\Theta)[\Delta]=\mathbb{E}\left<\Big[\frac{\partial\mathcal{L}}{\partial Z},D\Big]\right>_{F}=\alpha\cdot\mathbb{E}\Bigg[\underbrace{\sum_{R_{i,j}=0}D_{i,j}}_{\begin{subarray}{c}\text{penalty from}\\ \text{cross component}\end{subarray}}-\underbrace{\sum_{R_{i,j}=1}\frac{1-\phi_{\epsilon}(Z_{i,j})}{\phi_{\epsilon}(Z_{i,j})}D_{i,j}}_{\begin{subarray}{c}\text{weighted reward on}\\ \text{under-predicted positives}\end{subarray}}\Bigg].$ (9)
    In particular, $D_{B_{\ell}}\mathcal{R}(\Theta)[\Delta]\geq 0$ iff cross-component penalty $\geq$ within-component reward. Throughout this section, we will be using these names to denote the two competing sums whenever an expression like equation 9 appears.
@@
    Let $\Delta=\lvert B_{\ell}\rvert$ (entrywise absolute value) and let $D=\frac{\partial Z}{\partial B_{\ell}}[\lvert B_{\ell}\rvert$]. If, with positive probability under $\mathsf{ER}(n,p)$, $\Delta$ activates at least one cross-component pair, and if the cross component penalty term strictly dominates the within-component reward, then $B_{\ell}=0$. Equivalently, under activation at $\Delta=\lvert B_{\ell}\rvert$ and strict dominance by cross-component penalty, the only KKT stationary point in the $J_{n}$-channel is $B_{\ell}=0$.
@@
For the second claim, evaluate equation 9 at $\Delta=\lvert B_{\ell}\rvert$. Under the activation premise (Section C.3) and strict dominance by cross-component penalty, we obtain $D_{B_{\ell}}\mathcal{R}(\Theta)[\lvert B_{\ell}\rvert]=\left<\nabla_{B_{\ell}}\mathcal{R}(\Theta),\Delta\right>_{F}>0$. Since $\nabla_{B_{\ell}}\mathcal{R}(\Theta)\geq 0$ and $\lvert B_{\ell}\rvert\geq 0$, a strictly positive inner product violates the KKT complementary condition $\nabla_{B_{\ell}}\mathcal{R}(\Theta)\odot B_{\ell}=0$ unless $B_{\ell}=0$. ∎
@@
While Theorem C.5 mostly discusses the suppression of $B_{\ell}$, its (i) in fact reveals a quite interesting, opposite phenomenon: early training promotes $B_{\ell}$. Before the model learns to pick up easy connected pairs, the corresponding values $\phi_{\epsilon}(Z_{i,j})\ll 1$. Consequently, the fractions $(1-\phi_{\epsilon}(Z_{i,j}))/\phi_{\epsilon}(Z_{i,j})$ are large, making equation 9 negative. Gradient descent then pushes $B_{\ell}$ up “without feeling pressure.” As training proceeds, these easy connected pairs saturate ($\phi_{\epsilon}(Z_{i,j})\to 1$), while simultaneously $\Delta$ begins to active cross pairs (the “moreover” part of Section C.3), increasing the $R=0$ term in equation 9 and potentially flipping the sign. This is when the $J$-channel starts to incur penalty. This explains the transient “Phase 1” in §4.3.
@@
Recall $W_{\ell}=A_{\ell}\otimes I_{n}+B_{\ell}\otimes J_{n}$ and Section C.3. The $I$-channel controls local propagation within components; the $J$-channel couples to the global / mean direction and injects dense rank-one terms. In this section, we first shift to a micro-level perspective, focusing on the effects of individual samples (graphs), and then draw connection to how the training distribution determines the model’s eventual behavior (algorithmic vs. heuristic, §4.3).
@@
    (Within capacity) If $\mathrm{diam}(G)\leq 3^{L}$, then $D_{A_{\ell}}\mathcal{L}_{G}(\Theta)[\Delta_{A}]\leq 0$, with strict <0<0 whenever $\Delta_{A}$ is active on at least one unsaturated within-component pair. If, in addition, $G$ is disconnected, then $D_{B_{\ell}}\mathcal{L}_{G}(\Theta)[\Delta_{B}]>0$ if both of the following hold: $\Delta_{B}$ is active at at least one cross-component pair, and cross component penalty dominates within-component reward (cf. equation 9).
@@
To prove this Theorem, we split the argument into the following four lemmas, each isolating one ingredient of the dynamics. Firstly, Section C.4 shows that the local $I$-channel is monotone: any nonnegative $A_{\ell}$ cannot increase the loss and is strictly helpful on unsaturated within-component pairs. This lets us treat local corrections as “harmless,” while Section C.4 analyze the sign of the global $J$-channel (connected vs. disconnected), and Section C.4 determines which pairs are ever affected when $B=0$. Together, they yield the two cases in Theorem C.9.
@@
with strict negativity whenever $\phi_{\epsilon}(Z_{i,j})<1$. Entrywise, nonnegativity of the forward map (Section C.3) gives $\frac{\partial Z}{\partial A_{\ell}}[\Delta]\geq 0$. Therefore the Frobenius inner product $\leq 0$, and <0<0 under the stated conditions. ∎
@@
    hence $D_{B_{\ell}}\mathcal{L}_{G}(\Theta)[\Delta]\geq 0$ whenever the cross component mass penalty dominates the ratio-weighted within-component reward. Strict >0>0 holds if the inequality is strict, and $\Delta$ is active on at least one cross pair.
@@
By Section C.3 , $D\geq 0$ entrywise; moreover, $D_{i,j}>0$ exactly on pairs where $\Delta$ is active.
@@
Recall from Section 4.2: for a depth $L$ and a graph $G$ with adjacency matrix $A$, we call a pair $(i,j)$ within capacity if $A^{3^{L}}]_{i,j}>0$ and beyond capacity otherwise.
@@
Since $B_{\ell}=0$ implies $h_{\ell-1}W_{\ell}h_{\ell-1}^{\top}=\sum_{p,q}(A_{\ell})_{p,q}X_{p}X_{q}^{\top}$ from Section C.3, it is easy to see that they are block-diagonal w.r.t. connected components. Hence Section B.2 applies and support expands by at most a factor of $3$ per layer, and only within-capacity pairs receive mass. The density statement follows from Section C.3: For any $\Delta\geq 0$ in the $J$-channel, we have $\frac{\partial Z}{\partial B_{\ell}}=\sum_{p,q}\Delta_{p,q}u_{p}u_{q}^{\top}u\geq 0$. The strict positiveness characterization follows directly from Section C.3. ∎
@@
    (Within capacity) By Section C.4, for any $\Delta_{A}\geq 0$ the $I$-channel directional derivative is $\leq 0$, with strict inequality under the stated conditions. The result on disconnected graphs $G$ follows from Section C.4.
@@
    By Section C.4, with $B=0$, only within-capacity pairs can be affected by the $I$-channel, so Section C.4 gives $D_{A_{\ell}}\mathcal{L}_{G}(\Theta)[\Delta_{A}]\leq 0$. Since $G$ is connected and $\mathrm{diam}(G)>3^{L}$, there will be unsaturated pairs; then Section C.4(i) yields $D_{B_{\ell}}\mathcal{L}_{G}(\Theta)[\Delta_{B}]<0$, as claimed. ∎
@@
    (Local) From Section C.4, once the global $J$-channel has been suppressed, the local $I$-channel is consistently promoted until saturation.
@@
[…]
@@
      If $G$ is disconnected and $\mathrm{diam}(G)\leq 3^{L}$, then by Section C.4, $D_{B_{\ell}}\,\mathcal{L}_{G}(\Theta)[\Delta_{B}]\ \geq\ 0$ with strict >0>\!0 if cross-component errors persist (the $\sum_{R=0}D$ term strictly dominates the $\sum_{R=1}(1-\phi_{\epsilon}(Z))/\phi_{\epsilon}(Z)\cdot D$ term), and if $\Delta_{B}$ is active on cross pairs (i.e. $D^{(B)}_{ij}>0$ for some $R_{ij}=0$). The latter holds by Section C.3 if $\Delta_{B}$ is active on at least one cross pair. Integrating thus yields

### v2

@@ C.1 Characterizing Block Weights $W_{\ell}$
To do so, the proof is split into two parts: we prove that each $n\times n$ block must be diagonal using (ii) and Lemma B.2, and that the diagonal entries must realize the said forms by examining the forward maps under a curated, parameterized class of initial hidden states.
Step 1: each block must be diagonal. In this step, we argue that if a block admits a positive off-diagonal entry, then the certificate trick from Lemma B.1 will create a false positive entry on some output, contradicting (ii).
@@
Consider feeding the graph $G_{0}$ of $n$ isolated vertices into $\mathsf{TF}_{\Theta}^{L}$, so that $h_{0}(G_{0})=[I_{n}\mid I_{n}]$. The premises of Lemma B.2 hold, so every column of every $h_{\ell}(G_{0})$ is supported in exactly one row, which we called its label in $\{1,\ldots,n\}$. Hence, if we write $h_{\ell}(G_{0})=[X_{1}^{(\ell)}\mid\ldots\mid X_{K_{\ell}}^{(\ell)}]$ of contiguous $n\times n$ blocks, then each such block $X_{r}^{(\ell)}$ must be nonnegative and diagonal. Now expand
@@
  • If both $(i,r)$ and $(k,s)$ are activatable, take graphs $G_{i},G_{k}$ that make $X_{r}^{(\ell-1)}(G_{i})[i,i]>0$ and $X_{x}^{(\ell-1)}(G_{k})[k,k]>0$. Using the certificate mechanism in Lemma B.1, each positiveness admits a finite certificate subgraph with at most $2\cdot 3^{\ell-1}$ vertices. We then create a new graph $G^{\prime}$ and disjointly embed both certificates into it, leaving all other vertex isolated. The two labels $i,k$, viewed as nodes, now lie in different components. But then the product
@@
Suppose an $L$-layer Disentangled Transformer $\mathsf{TF}_{\Theta}^{L}$ has nonnegative parameters. Suppose $\mathsf{TF}_{\Theta}^{L}$ is layerwise permutation equivariant, i.e., for each $\ell$, any hidden states $h\in\mathbb{R}^{n\times d_{\ell-1}}$, and any permutation $P\in S_{n}$,
@@ C.2 Population Gradient Lives in the Equivariant Algebra
Under 4.8, in particular using layerwise parameterization $W_{\ell}=A_{\ell}\otimes I_{n}+B_{\ell}\otimes J_{n}$, fix a layer $\ell$ and let $K=K_{\ell-1}$. Then the population gradient with respect to $W_{\ell}$ lies in $M_{K}(\mathbb{R})\otimes\mathrm{span}\{I_{n},J_{n}\}$: there exist matrices $G_{\ell}^{(I)},G_{\ell}^{(J)}\in\mathbb{R}^{K\times K}$ such that
@@ C.3 Which Conditions Encourage $W_{\ell}\approx A_{\ell}\otimes I_{n}$?
Throughout the analysis of training dynamics, we inherit the notations used in 4.8: we use $Z$ to denote the model output, $R$ the reachability matrix, $A$ the adjacency matrix, $\mathcal{L}=\mathcal{L}(Z;R)$ the loss, and $\mathcal{R}(\Theta)$ the population risk $\mathcal{R}(\Theta):=\mathbb{E}_{G\sim\mathrm{ER}(n,p)}[\mathcal{L}(\mathsf{TF}_{\Theta}^{L}(A_{G});R_{G})]$.
@@
Assume 4.8. Fix any layer $\ell$ and decompose $W_{\ell}=A_{\ell}\otimes I_{n}+B_{\ell}\otimes J_{n}$. Let $Z$ be the output, $R$ the reachability matrix (ground truth), $\mathcal{L}=\mathcal{L}(Z;R)$ the loss, and $\mathcal{R}(\Theta)$ the population risk.
@@
    $D_{B_{\ell}}\mathcal{R}(\Theta)[\Delta]=\mathbb{E}\left<\Big[\frac{\partial\mathcal{L}}{\partial Z},D\Big]\right>_{F}=\alpha\cdot\mathbb{E}\Bigg[\underbrace{\sum_{R_{i,j}=0}D_{i,j}}_{\begin{subarray}{c}\text{cross component}\\ \text{penalty}\end{subarray}}-\underbrace{\sum_{R_{i,j}=1}\frac{1-\phi_{\epsilon}(Z_{i,j})}{\phi_{\epsilon}(Z_{i,j})}D_{i,j}}_{\begin{subarray}{c}\text{within-component}\\ \text{reward}\end{subarray}}\Bigg].$ (6)
    In particular, $D_{B_{\ell}}\mathcal{R}(\Theta)[\Delta]\geq 0$ iff the Population-Level Dominance Condition holds (i.e. Equation 6 is positive). Throughout this Appendix, we will use “cross component penalty” and “within-component reward” to denote these two competing terms.
@@
    Let $\Delta=\lvert B_{\ell}\rvert$ (entrywise absolute value) and let $D=\frac{\partial Z}{\partial B_{\ell}}[\lvert B_{\ell}\rvert$]. If, with positive probability under $\mathsf{ER}(n,p)$, $\Delta$ activates at least one cross-component pair, and if the Population-Level Dominance Condition holds, then $B_{\ell}=0$. Equivalently, under activation at $\Delta=\lvert B_{\ell}\rvert$ and strict dominance by cross-component penalty, the only KKT stationary point in the $J_{n}$-channel is $B_{\ell}=0$.
@@
For the second claim, evaluate equation 6 at $\Delta=\lvert B_{\ell}\rvert$. Under the activation premise (Lemma C.6) and strict dominance by cross-component penalty, we obtain $D_{B_{\ell}}\mathcal{R}(\Theta)[\lvert B_{\ell}\rvert]=\left<\nabla_{B_{\ell}}\mathcal{R}(\Theta),\Delta\right>_{F}>0$. Since $\nabla_{B_{\ell}}\mathcal{R}(\Theta)\geq 0$ and $\lvert B_{\ell}\rvert\geq 0$, a strictly positive inner product violates the KKT complementary condition $\nabla_{B_{\ell}}\mathcal{R}(\Theta)\odot B_{\ell}=0$ unless $B_{\ell}=0$. ∎
@@
While Theorem C.5 mostly discusses the suppression of $B_{\ell}$, its (i) in fact reveals a quite interesting, opposite phenomenon: early training promotes $B_{\ell}$. Before the model learns to pick up easy connected pairs, the corresponding values $\phi_{\epsilon}(Z_{i,j})\ll 1$. Consequently, the fractions $(1-\phi_{\epsilon}(Z_{i,j}))/\phi_{\epsilon}(Z_{i,j})$ are large, making equation 6 negative. Gradient descent then pushes $B_{\ell}$ up “without feeling pressure.” As training proceeds, these easy connected pairs saturate ($\phi_{\epsilon}(Z_{i,j})\to 1$), while simultaneously $\Delta$ begins to active cross pairs (the “moreover” part of Lemma C.6), increasing the $R=0$ term in equation 6 and potentially flipping the sign. This is when the $J$-channel starts to incur penalty. This explains the transient “Phase 1” in §4.3.
@@
Recall $W_{\ell}=A_{\ell}\otimes I_{n}+B_{\ell}\otimes J_{n}$ and Lemma C.6. The $I$-channel controls local propagation within components; the $J$-channel couples to the global / mean direction and injects dense rank-one terms. In this section, we first shift to a micro-level perspective, focusing on the effects of individual samples (graphs), and then draw connection to how the training distribution determines the model’s eventual behavior (algorithmic vs. heuristic, §4.3).
@@
    (Within capacity) If $\mathrm{diam}(G)\leq 3^{L}$, then $D_{A_{\ell}}\mathcal{L}_{G}(\Theta)[\Delta_{A}]\leq 0$, with strict <0<0 whenever $\Delta_{A}$ is active on at least one unsaturated within-component pair. If, in addition, $G$ is disconnected, then $D_{B_{\ell}}\mathcal{L}_{G}(\Theta)[\Delta_{B}]>0$ if both of the following hold: $\Delta_{B}$ is active at at least one cross-component pair, and Population-Level Dominance Condition holds.
@@
To prove this Theorem, we split the argument into the following four lemmas, each isolating one ingredient of the dynamics. Firstly, Lemma C.10 shows that the local $I$-channel is monotone: any nonnegative $A_{\ell}$ cannot increase the loss and is strictly helpful on unsaturated within-component pairs. This lets us treat local corrections as “harmless,” while Lemma C.11 analyze the sign of the global $J$-channel (connected vs. disconnected), and Lemma C.12 determines which pairs are ever affected when $B=0$. Together, they yield the two cases in Theorem C.9.
@@
with strict negativity whenever $\phi_{\epsilon}(Z_{i,j})<1$. Entrywise, nonnegativity of the forward map (Lemma C.6) gives $\frac{\partial Z}{\partial A_{\ell}}[\Delta]\geq 0$. Therefore the Frobenius inner product $\leq 0$, and <0<0 under the stated conditions. ∎
@@
    hence $D_{B_{\ell}}\mathcal{L}_{G}(\Theta)[\Delta]\geq 0$ whenever the Population-Level Dominance Condition holds. Strict >0>0 holds if the inequality is strict, and $\Delta$ is active on at least one cross pair.
@@
By Lemma C.6 , $D\geq 0$ entrywise; moreover, $D_{i,j}>0$ exactly on pairs where $\Delta$ is active.
@@
Recall from Definition 4.6: for a depth $L$ and a graph $G$ with adjacency matrix $A$, we call a pair $(i,j)$ within capacity if $A^{3^{L}}]_{i,j}>0$ and beyond capacity otherwise.
@@
Since $B_{\ell}=0$ implies $h_{\ell-1}W_{\ell}h_{\ell-1}^{\top}=\sum_{p,q}(A_{\ell})_{p,q}X_{p}X_{q}^{\top}$ from Lemma C.6, it is easy to see that they are block-diagonal w.r.t. connected components. Hence Lemma B.2 applies and support expands by at most a factor of $3$ per layer, and only within-capacity pairs receive mass. The density statement follows from Lemma C.6: For any $\Delta\geq 0$ in the $J$-channel, we have $\frac{\partial Z}{\partial B_{\ell}}=\sum_{p,q}\Delta_{p,q}u_{p}u_{q}^{\top}u\geq 0$. The strict positiveness characterization follows directly from Lemma C.6. ∎
@@
    (Within capacity) By Lemma C.10, for any $\Delta_{A}\geq 0$ the $I$-channel directional derivative is $\leq 0$, with strict inequality under the stated conditions. The result on disconnected graphs $G$ follows from Lemma C.11.
@@
    By Lemma C.12, with $B=0$, only within-capacity pairs can be affected by the $I$-channel, so Lemma C.10 gives $D_{A_{\ell}}\mathcal{L}_{G}(\Theta)[\Delta_{A}]\leq 0$. Since $G$ is connected and $\mathrm{diam}(G)>3^{L}$, there will be unsaturated pairs; then Lemma C.11(i) yields $D_{B_{\ell}}\mathcal{L}_{G}(\Theta)[\Delta_{B}]<0$, as claimed. ∎
@@
    (Local) From Lemma C.10, once the global $J$-channel has been suppressed, the local $I$-channel is consistently promoted until saturation.
@@
      If $G$ is connected and $\mathrm{diam}(G)>3^{L}$, then by Lemma C.12, every beyond-capacity pair has $Z_{ij}=0$ while $R_{ij}=1$. For those pairs, we have $\partial\mathcal{L}/\partial Z=-\alpha(1-\phi_{\epsilon}(Z))/\phi_{\epsilon}(Z)<0$. By Lemma C.6, the inner product $\left<\partial\mathcal{L}/\partial Z,\partial Z/\partial B_{\ell}[\Delta_{B}]\right>_{F}<0$ too. Integrating over all beyond-capacity, connected graphs yields
@@
      If $G$ is disconnected and $\mathrm{diam}(G)\leq 3^{L}$, then by Lemma C.11, $D_{B_{\ell}}\,\mathcal{L}_{G}(\Theta)[\Delta_{B}]\ \geq\ 0$ with strict >0>\!0 if cross-component errors persist (the $\sum_{R=0}D$ term strictly dominates the $\sum_{R=1}(1-\phi_{\epsilon}(Z))/\phi_{\epsilon}(Z)\cdot D$ term), and if $\Delta_{B}$ is active on cross pairs (i.e. $D^{(B)}_{ij}>0$ for some $R_{ij}=0$). The latter holds by Lemma C.6 if $\Delta_{B}$ is active on at least one cross pair. Integrating thus yields